Isolation Method of Rate Law Determination

Consider the base hydrolysis of an ester RCOOR + OH-→ RCOO-+ ROH (where R and Rrepresent organic groups) The kinetics of this reaction, for a particular ester, were studied using the isolation method When [OH>> [RCOOR] a pseudo-rate law can be posed: With initial concentrations: [OH-]0 = 0.800 M [RCOOR]0 = 0.0010 M, a plot of In[RCOORT versus t gave a straight line with slope =-1.01 s-1 and y-intercept =-6.91 What is the order n? What is the corresponding pseudo rate constant k? (include appropriate units) k 1.01 /s pts You are correct. Previous Tries When [OH ]<<[RCOOR] a pseudo-rate law can be posed: With initial concentrations: [OH-]0 = 0.0015 M [RCOORJo = 0.700 M a plot of In[OH-] versus t gave a straight line with slope =-0.881 s-1 and y-intercept =-6.50 What is the order m? What is the corresponding pseudo rate constant k? (include appropriate units) pts You are correct. Previous Tries k= 0.88 1 /s The overall rate law can be posed as: d RCOOR Wha nstant, k, for the reaction? (include appropriate units) pts Submit Answer | Incompatible units. No conversion found between 1/s and the required units. Tries 0/5 Previous Tries
Answer #1

for the 1st case, when [OH]>>[RCOOR']

K= 1.01/s , rate law =K' [RCOOR']

K'= K[OH]

K= K'/[OH] = 1.01/0.8= 1.26/M.sec

for the second case, K''= K'[RCOOR]

K' = K''/[RCOOR]= 0.881/0.70=1.26 /M.sec

This is the rate constant for the overalll reaction